9x-2y=15
4x+3y=-5
Solve by elimination

Answer by checkley79(3341)

9x-2y=15 MULTIPLY BY 3
4x+3y=-5 MULTIPLY BY 2 & ADD.
27X-6Y=45
8X+6Y=-10
--------------
35X=35
X=35/35
X=1 ANS.
9*1-2Y=15
9-2Y=15
-2Y=15-9
-2Y=6
Y=6/-2
Y=-3 ANS.
PROOF:
4*1+3*-3=-5
4-9=-5
-5=-5